OpenStudy (anonymous):

A survey of 260 families showed that 99 had a dog; 76 had a cat; 34 had a dog and a cat; 98 had neither a cat nor a dog nor a parakeet; 8 had a cat and a dog and a parakeet. How many had a parakeet only?

OpenStudy (anonymous):

 Now, Dog+Cat=99+76-34 (by set theory and Venn diagram now, notice that Dog+Cat+X=260-98. Do you need further help?

OpenStudy (anonymous):

(76 - 34) + (99 - 34) + 34 + P = 162

OpenStudy (anonymous):

these are my choices A 21 B 36 C 31 D 26

OpenStudy (anonymous):

(76 - 34) + (99 - 34) + 34 + P = 162 Can you solve for p ?

OpenStudy (anonymous):

42+65+34+p=162?

OpenStudy (anonymous):

 Do you understand this ?

OpenStudy (anonymous):

i got 21

OpenStudy (anonymous):

A survey of 86 college students was taken to determine the musical styles they liked. Of those, 36 students listened to rock, 34 to classical, and 37 to jazz. Also, 17 students listened to rock and jazz, 15 to rock and classical, and 16 to classical and jazz. Finally, 6 students listened to all three musical styles. Construct a Venn diagram and determine the cardinality for each region in order to answer the following question. How many students listened to classical and jazz, but not rock?